OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any.-3x-4y=2, 8y=-6x-4

OpenStudy (anonymous):

On the first one you would add -3x to both sides, then divide both sides by -4

OpenStudy (anonymous):

The second one you divide both sides by 8

OpenStudy (anonymous):

You have to use solve by substitution

jhonyy9 (jhonyy9):

8y= -6x-4 8y=2(-3x-2) divide both sides by 2 4y=-3x-2 2=-3x-4y so then you check it ,this equation is the same like first